How many moles of air are in the lungs of a average person with a total lung capacity of 3.8 L? Assume that the person is at 1.0 atm pressure and has a normal body temperature of 37 C
Answer #1

This problem is based on Ideal gas law. Assuming air in lungs behave as an ideal gas we can determine the number of moles.

PV = nRT

n = PV / RT...where P = 1.0atm, V = 3.8L,  T = 37oC = 37+273 = 310 K

(1.00 atm)(3.8 L) - = 0.15 mol (0.08206 kuam. )(310 K)

n = (1*3.8) / (0.08206*310) = 0.15 mole

Thus the number of moles of air in the lungs is = 0.15 mole